Our Blog

System Status news and updates

Router Upgrade (2011-10-25 00:00-06:00)

Posted by Support Team on 18 10 2011. 0 Comments

Our upstream provider will be performing updates to a number of routers between 00:00 and 06:00 on the 25th October 2011. During this period we anticipate a period of 30 minutes that will be ‘at risk’, within this window you may experience intermittent issues accessing services. This activity does not…

VLINWEB1 file system (2011-08-05)

Posted by Support Team on 05 08 2011. 0 Comments

We experienced file system corruption on Linux server VLINWEB1 which caused inconsistent serving of web pages and email. A ‘fskc’ has been completed and the server rebooted. All services on this server are now operational. If you are still experiencing an issue please contact support and we will investigate your specific…

Core Router Reboot (2011-07-28)

Posted by Support Team on 25 07 2011. 0 Comments

Our data centre provider has advised this morning that they will be performing maintenance on 28th July between 00:05 and 02:00. This will involve them updating the firmware on one of their routers which we anticipate will result in a short outage of 5 to 15 minute during this window. During…

POP3 mail issue [resolved]

Posted by Support Team on 11 03 2011. 0 Comments

We experienced an issue with users attempting POP3 based connections to our mail servers for approximately 1 hour. We have restored service and now investigating the root cause. During this issue customers connecting using IMAP or Webmail were unaffected.…

Packet Loss

Posted by Support Team on 02 01 2011. 2 Comments on Packet Loss

Our monitoring has highlighted a packet loss issue which is impacting our external connectivity. Customers may see this as slow or intermittent access to services. We are currently in contact with our upstream network provider.…

Cluster Node Issue

Posted by Support Team on 07 10 2010. 0 Comments

We experienced a network issue this evening on one of our cluster nodes this evening at approximately 23:00 which resulted in some websites and email being unavailable for approximately 10 minutes. Further investigations will be conducted to isolate the cause of the crash and the reason why automated recovery did not occur.…

Connectivity issue to some locations

Posted by Support Team on 06 10 2010. 1 Comment on Connectivity issue to some locations

Our monitoring has identified an issue exchanging traffic with some third party networks. We are currently investigationg this issue.…

Generator Test (2010-09-11)

Posted by Support Team on 10 09 2010. 1 Comment on Generator Test (2010-09-11)

A data centre generator test will occur at 09:30 tomorrow morning. There should be no impact to services as UPS protection will power devices whilst the data centre switch from mains to generators and back again.…

System Outage

Posted by Support Team on 13 03 2010. 1 Comment on System Outage

This morning we experienced a system outage due to an issue at our partner’s data centre. We are currently investigating the issue with them and will provide further specific information once we have it. During the outage our offsite DNS and Mail services ensures that your domain names continue to resolve…

MySQL Upgrade

Posted by Support Team on 28 09 2009. 0 Comments

This evening we have implemented a number of upgrades to the Windows hosting MySQL which will offer much improved performance. A combination of hardware upgrades and configuration optimisation have resulted a visible increase in performance.…